# Whom do Explanations Serve? A Systematic Literature Survey of User Characteristics in Explainable Recommender Systems Evaluation

### v1

@@ 1. Introduction
Recommender systems have become part of many people’s everyday online interactions. Whether it is suggestions for what movie to watch, what items to buy, which news to read, or which restaurant to visit, these systems generally reach and serve a broad audience in their decision-making process. Adding explanations to recommender systems is said to serve a multitude of benefits. First, by providing the reasoning behind specific recommendations, explanations can mitigate issues dealing with users’ autonomy that can occur when recommendations nudge or persuade users in a particular direction that is not beneficial for them (Milano et al., 2020). Carefully designed explanations can help the user to identify when the system is “wrong” and provide options to correct it (see the scrutability goal of Tintarev and Masthoff (2007)). Second, they can help users understand why they receive a specific recommendation, increase their trust, or allow them to make better decisions (Tintarev and Masthoff, 2007; Vultureanu‐Albişi and Bădică, 2022). The quality of recommender systems explanations is often assessed based on whether it has such effects on a user. Measuring the explanation quality with offline metrics is still an open research challenge (Zhang et al., 2020). Text-based explanations have been evaluated with metrics from the natural language processing domain. Ariza-Casabona et al. (2024), for example, use BLEU-n, ROUGE-n, and BERT-S along with content repetition, explanation length, and sentence uniqueness measures to measure and compare the text quality of the generated explanations. These metrics are, however, limited to text-based explanations. Other metrics include ranking- and retrieval-based metrics, such as Recall and NDCG used in Yu et al. (2024). Both metric types have the downside that they require a ground truth to compare the generated explanations, which is often approximated through online reviews. The results from using online reviews as a ground truth need to be carefully discussed and interpreted as they can be biased (Hu et al., 2017) and might exclude certain sociodemographics and personality types (Manner and Lane, 2017). Liao et al. (2020) further point out that the usefulness of explanations for users is dependent on their motivation for using the system and their downstream actions, which cannot be captured by current offline evaluation metrics.
@@
In the first research question, we look at the description of the participants in the user studies. We analyze the participants’ demographic information, personality traits, and prior experience to provide more context to the evaluated explanation effects. Subsequently, we discuss what user groups have not been included when evaluating recommender systems explanations based on the results of RQ 1. By creating awareness about these gaps, we hope to generate opportunities to specifically guide research endeavors in a direction that has improved coverage of the entire range of recommender system users. For the second research question, we look at the findings regarding the explanations’ effects on users and investigate how these might differ for specific user characteristics. We look at and discuss the papers to show possible indications for the effects that have been evaluated and highlight gaps where more research is needed.
@@ 3.1. Data Collection and Selection
The selection process narrowed the corpus to 124 papers containing a total of 158 user studies (some papers reported on multiple studies), from which we extracted the information reported on the user study participants and the results. [^9] The full list of papers and categorization can be viewed here: https://docs.google.com/spreadsheets/d/1v1u9RVNEmnZ2wROmySizBaYyFcMgRGn0 In Section 3.2, we elaborate on the annotation process of the papers in our corpus.
@@ 4.2. The Impact of User Characteristics
Table 4. Number of results and publications in which the effect of user characteristics on explanation effects was analyzed. Rows with ✓indicate that an impact of the user characteristic was found; rows with $\times$ indicate that no impact was found. The last column sums the number of findings for each user characteristic. The last row sums up the findings where an effect was found or not found. The superscript letter indicates the explanation type: $t$: textual, $v$: visual, $h$: hybrid, $s$: schematic (i.e., a table), and $n$: numerical. Findings can contain multiple explanations of different types that were grouped together. Subscript letters show the explanation method: $p$: post-hoc, $i$: intrinsic. Application domains are depicted as follows: [Uncaptioned image]: document, [Uncaptioned image]: e-commerce, [Uncaptioned image]: education, [Uncaptioned image]: energy saving, [Uncaptioned image]: health, [Uncaptioned image]: movie, [Uncaptioned image]: music, [Uncaptioned image]: online dating, [Uncaptioned image]: point-of-interest, [Uncaptioned image]: social. The recommender type is shown after the domain and indicated by cbcb for content-based, cfcf for collaborative filtering, and hyhy for hybrid systems. Articles where information on any of the categories was not provided or the recommendations were simulated are marked with -.
@@
Table 4 shows an overview of the evaluated explanation effects, user characteristics, the evaluated explanation type, explanation method, type of recommender system, and the application domain in which the explanation was evaluated. At first glance, the data is sparse for all combinations of effects and characteristics, making it impossible to determine a statistically significant impact of user characteristics on explanation effects. We can identify some indications that might lead to the generation of hypotheses that can be tested in further experiments.
@@ 4.2.4. Explanation Context
We looked at four additional variables that provide context to the results: explanation type, explanation generation method, recommender type, and the application domain in which the explanation was evaluated. In terms of explanation type, we found 68 results in which multiple explanations of different types were combined in the result. We cannot determine any patterns in instances where the result addresses only one explanation. We often see that an impact of a user characteristic was found and not found for the same explanation type (see the trust or usability/UX effects for the trust propensity user characteristic in Table 4 for an example), but no explanation type seems to be more likely to evoke an effect. The explanation generation method is predominantly either unknown or the explanations were simulated for the user evaluation. 103 out of 165 findings are based on these explanations. Intrinsically explainable recommender systems are rare in this evaluation. Only 21 of the findings stem from a model intrinsic explanation such that a comparison between intrinsic and post-hoc explanations is not possible. We identified a total of nine different application domains that were evaluated, with point-of-interest recommendations being the most common (44 findings), followed by document (39 findings), and social recommendations (27 findings). We notice the dominance of some domains in certain user characteristics. The decision-making type and social awareness, for example, were almost exclusively evaluated in the point-of-interest domain and trust propensity and the education level of participants in the document and education domain, respectively. The dominant type of recommender system is content-based, with 107 out of 165 findings. Pure collaborative filtering systems are only used to generate the recommendations for 10 findings.

### v2

@@ 1. Introduction
Recommender systems have become part of many people’s everyday online interactions. Whether it is suggestions for what movie to watch, what items to buy, which news to read, or which restaurant to visit, these systems generally reach and serve a broad audience in their decision-making process. Adding explanations to recommender systems is said to serve a multitude of benefits. First, by providing the reasoning behind specific recommendations, explanations can mitigate issues dealing with users’ autonomy that can occur when recommendations nudge or persuade users in a particular direction that is not beneficial for them (Milano et al., 2020). Carefully designed explanations can help the user to identify when the system is “wrong” and provide options to correct it (see the scrutability goal of Tintarev and Masthoff (2007)). Second, they can help users understand why they receive a specific recommendation, increase their trust, or allow them to make better decisions (Tintarev and Masthoff, 2007; Vultureanu‐Albişi and Bădică, 2022). The quality of recommender systems explanations is often assessed based on whether it has such effects on a user. Measuring the explanation quality with offline metrics is still an open research challenge (Zhang et al., 2020). Text-based explanations have been evaluated with metrics from the natural language processing domain. Ariza-Casabona et al. (2024), for example, use BLEU-n, ROUGE-n, and BERT-S along with content repetition, explanation length, and sentence uniqueness measures to measure and compare the text quality of the generated explanations. In addition to BLEU and ROUGE, Kokubo and Sugiyama (2022) and Yu et al. (2023) evaluate the diversity of the generated sentences in the explanation with the Distinct score. Yu et al. (2023) additionally use METEOR, a score based on unigram matching between the generated text and a ground truth. These metrics are, however, limited to text-based explanations. Other metrics include ranking- and retrieval-based metrics, such as Recall and NDCG used in Yu et al. (2024). Except for the diversity-based metrics, these metric types have the downside that they require a ground truth to compare the generated explanations, which is often approximated through online reviews. The results from using online reviews as a ground truth need to be carefully discussed and interpreted as they can be biased (Hu et al., 2017) and might exclude certain sociodemographics and personality types (Manner and Lane, 2017). Liao et al. (2020) further point out that the usefulness of explanations for users is dependent on their motivation for using the system and their downstream actions, which cannot be captured by current offline evaluation metrics.
@@
In the first research question, we look at the description of the participants in the user studies. We analyze the participants’ demographic information, personality traits, and prior experience to provide more context to the evaluated explanation effects. Subsequently, we discuss what user groups have not been included when evaluating recommender systems explanations based on the results of RQ 1. By creating awareness about these gaps, we hope to generate opportunities to specifically guide research endeavors in a direction that has improved coverage of the entire range of recommender system users. For the second research question, we look at the findings regarding the explanations’ effects on users and investigate how these might differ for specific user characteristics. We look at and discuss the papers to show possible indications for the effects that have been evaluated and highlight gaps where more research is needed. [^3] We provide an interactive version of this analysis here: https://kathriwa.github.io/interactive-survey-visualization/#/.
@@ 3.1. Data Collection and Selection
The selection process narrowed the corpus to 124 papers containing a total of 158 user studies (some papers reported on multiple studies), from which we extracted the information reported on the user study participants and the results. [^10] The full list of papers and categorization can be viewed here: https://doi.org/10.5281/zenodo.14771123 In Section 3.2, we elaborate on the annotation process of the papers in our corpus.
@@ 4.2. The Impact of User Characteristics
Table 4. Number of results and publications in which the effect of user characteristics on explanation effects was analyzed with the explanation type and explanation generation method as context variables. Rows with ✓indicate that an impact of the user characteristic was found; rows with $\times$ indicate that no impact was found. The last column sums the number of findings for each user characteristic. The last row sums up the findings where an effect was found or not found. The superscript letter indicates the explanation type: $t$: textual, $v$: visual, $h$: hybrid, $s$: schematic (i.e., a table), and $n$: numerical. Findings can contain multiple explanations of different types that were grouped together. Subscript letters show the explanation method: $p$: post-hoc, $i$: intrinsic. Articles where information on any of the categories was not provided or the recommendations were simulated are marked with -.
@@
We present these results in Tables 4, 5, and 6 with different context variables [^11] An interactive version of these tables is available at https://kathriwa.github.io/interactive-survey-visualization/#/characteristics. More precisely, we present the evaluated explanation effects on different user characteristics using as context variable(s) the evaluated explanation type and explanation method in Table 4, the application domain in Table 5, and the type of recommender systems in Table 6. At first glance, the data is sparse for all combinations of effects and characteristics, making it impossible to determine a statistically significant impact of user characteristics on explanation effects. We can identify some indications that might lead to the generation of hypotheses that can be tested in further experiments.
@@ 4.2.4. Explanation Context
We looked at four additional variables that provide context to the results: explanation type, explanation generation method, recommender type, and the application domain in which the explanation was evaluated. In terms of explanation type, we found 68 results in which multiple explanations of different types were combined in the result. We cannot determine any patterns in instances where the result addresses only one explanation. We often see that an impact of a user characteristic was found and not found for the same explanation type (see the trust or usability/UX effects for the trust propensity user characteristic in Table 4 for an example), but no explanation type seems to be more likely to evoke an effect. The explanation generation method is predominantly either unknown or the explanations were simulated for the user evaluation. 103 out of 165 findings are based on these explanations. Intrinsically explainable recommender systems are rare in this evaluation. Only 21 of the findings stem from a model intrinsic explanation such that a comparison between intrinsic and post-hoc explanations is not possible.
Table 5. Number of results and publications in which the effect of user characteristics on explanation effects was analyzed with the application domain as a context variable. Rows with ✓indicate that an impact of the user characteristic was found; rows with $\times$ indicate that no impact was found. The last column sums the number of findings for each user characteristic. The last row sums up the findings where an effect was found or not found. Application domains are depicted as follows: [Uncaptioned image]: document, [Uncaptioned image]: e-commerce, [Uncaptioned image]: education, [Uncaptioned image]: energy saving, [Uncaptioned image]: health, [Uncaptioned image]: movie, [Uncaptioned image]: music, [Uncaptioned image]: online dating, [Uncaptioned image]: point-of-interest, [Uncaptioned image]: social.
The first two columns list the user characteristics that were analyzed in the evaluations of our corpus sorted by their category (demographic, personality, experience). Each characteristic consists of two rows, one in which the papers that found an effect of the user characteristic are referenced and one in which the papers that did not find an effect are referenced.
User
Characteristic
Effect
Found
Effectiveness Efficiency
Perceived Explanation
Quality
Persuasiveness Transparency Trust Usability/UX Total
Demographic Age (Wilkinson et al., 2021) [Uncaptioned image] (Nelekar et al., 2021) [Uncaptioned image] 2
$\times$
Location (Berkovsky et al., 2018) [Uncaptioned image] (Berkovsky et al., 2018) [Uncaptioned image] 6
$\times$
Gender (Kleinerman et al., 2018) [Uncaptioned image] (Ma et al., 2021) [Uncaptioned image] (Ma et al., 2021) [Uncaptioned image] (Ma et al., 2021) [Uncaptioned image], (Tran et al., 2020) [Uncaptioned image] 9
$\times$ (Wilkinson et al., 2021) [Uncaptioned image], (Coba et al., 2019) [Uncaptioned image] (Nelekar et al., 2021) [Uncaptioned image]
Education (Zheng and Toribio, 2021) [Uncaptioned image] (Ma et al., 2021) [Uncaptioned image] 6
$\times$ (Ma et al., 2021; Zheng and Toribio, 2021) [Uncaptioned image] (Ma et al., 2021) [Uncaptioned image]
Personality Agreeableness (Najafian et al., 2021a) [Uncaptioned image] 4
$\times$ (Najafian et al., 2021b) [Uncaptioned image] (Abdulrahman et al., 2022) [Uncaptioned image]
Conscientious- ness (Kouki et al., 2020) [Uncaptioned image] (Nelekar et al., 2021; Abdulrahman et al., 2022) [Uncaptioned image] 4
$\times$ (Abdulrahman et al., 2022) [Uncaptioned image]
Decision-Making Type (Hernandez-Bocanegra and Ziegler, 2021, 2020) [Uncaptioned image] (Hernandez-Bocanegra and Ziegler, 2020) [Uncaptioned image] (Hernandez-Bocanegra and Ziegler, 2021, 2020; Coba et al., 2019) [Uncaptioned image] (Hernandez-Bocanegra and Ziegler, 2021) [Uncaptioned image] (Hernandez-Bocanegra and Ziegler, 2021, 2020) [Uncaptioned image] (Hernandez-Bocanegra and Ziegler, 2021) [Uncaptioned image] 17
$\times$ (Coba et al., 2019) [Uncaptioned image] (Hernandez-Bocanegra and Ziegler, 2020) [Uncaptioned image]
Extraversion (Nelekar et al., 2021) [Uncaptioned image] 4
$\times$ (Najafian et al., 2021b) [Uncaptioned image] (Abdulrahman et al., 2022) [Uncaptioned image]
Need for Cognition (Martijn et al., 2022) [Uncaptioned image] (Gajos and Mamykina, 2022) [Uncaptioned image] 17
$\times$ (Chatti et al., 2022) [Uncaptioned image], (Guesmi et al., 2022) [Uncaptioned image] (Chatti et al., 2022) [Uncaptioned image], (Guesmi et al., 2022) [Uncaptioned image] (Liao and Sundar, 2021) [Uncaptioned image], (Millecamp et al., 2020) [Uncaptioned image] (Chatti et al., 2022) [Uncaptioned image], (Guesmi et al., 2022) [Uncaptioned image] (Gajos and Mamykina, 2022) [Uncaptioned image], (Chatti et al., 2022) [Uncaptioned image], (Guesmi et al., 2022) [Uncaptioned image] (Chatti et al., 2022) [Uncaptioned image], (Guesmi et al., 2022) [Uncaptioned image] (Chatti et al., 2022) [Uncaptioned image], (Guesmi et al., 2022) [Uncaptioned image]
Neuroticism (Nelekar et al., 2021) [Uncaptioned image],(Kouki et al., 2020) [Uncaptioned image] 7
$\times$ (Najafian et al., 2021a, b) [Uncaptioned image] (Abdulrahman et al., 2022) [Uncaptioned image]
Openness (Kouki et al., 2020) [Uncaptioned image] 3
$\times$ (Abdulrahman et al., 2022) [Uncaptioned image]
Personal Innovativeness (Chatti et al., 2022) [Uncaptioned image] (Chatti et al., 2022) [Uncaptioned image] (Guesmi et al., 2022) [Uncaptioned image] (Chatti et al., 2022) [Uncaptioned image] 13
$\times$ (Guesmi et al., 2022) [Uncaptioned image] (Guesmi et al., 2022) [Uncaptioned image] (Chatti et al., 2022) [Uncaptioned image], (Guesmi et al., 2022) [Uncaptioned image] (Chatti et al., 2022) [Uncaptioned image] (Chatti et al., 2022) [Uncaptioned image], (Guesmi et al., 2022) [Uncaptioned image] (Chatti et al., 2022) [Uncaptioned image], (Guesmi et al., 2022) [Uncaptioned image]
Trust Propensity (Nelekar et al., 2021) [Uncaptioned image] (Guesmi et al., 2022) [Uncaptioned image] 8
$\times$ (Guesmi et al., 2022) [Uncaptioned image] (Guesmi et al., 2022) [Uncaptioned image] (Guesmi et al., 2022) [Uncaptioned image] (Guesmi et al., 2022) [Uncaptioned image] (Guesmi et al., 2022) [Uncaptioned image] (Guesmi et al., 2022) [Uncaptioned image]
Rationality 3
$\times$ (Hernandez-Bocanegra et al., 2020) [Uncaptioned image] (Naveed et al., 2018) [Uncaptioned image]
Social Awareness (Hernandez-Bocanegra and Ziegler, 2021, 2020) [Uncaptioned image] (Hernandez-Bocanegra and Ziegler, 2021, 2020) [Uncaptioned image] (Hernandez-Bocanegra and Ziegler, 2021) [Uncaptioned image] (Starke et al., 2021b) [Uncaptioned image], (Hernandez-Bocanegra and Ziegler, 2021, 2020) [Uncaptioned image] (Hernandez-Bocanegra et al., 2020; Hernandez-Bocanegra and Ziegler, 2021, 2020) [Uncaptioned image] 15
$\times$ (Hernandez-Bocanegra and Ziegler, 2020) [Uncaptioned image] (Hernandez-Bocanegra et al., 2020) [Uncaptioned image]
Trust in Technology (Loepp et al., 2019) [Uncaptioned image] 1
$\times$
Valence (Kim et al., 2020) [Uncaptioned image] 1
$\times$
Experience Domain Experience (Loepp et al., 2019) [Uncaptioned image] 11
$\times$ (Guesmi et al., 2022) [Uncaptioned image] (Guesmi et al., 2022) [Uncaptioned image] (Kouki et al., 2020) [Uncaptioned image] (Guesmi et al., 2022) [Uncaptioned image], (Kouki et al., 2020) [Uncaptioned image] (Guesmi et al., 2022) [Uncaptioned image] (Guesmi et al., 2022) [Uncaptioned image] (Guesmi et al., 2022) [Uncaptioned image]
Technical Expertise (Jacobs et al., 2021) [Uncaptioned image] (Chatti et al., 2022) [Uncaptioned image] (Kouki et al., 2020) [Uncaptioned image] 16
$\times$ (Chatti et al., 2022) [Uncaptioned image], (Guesmi et al., 2022) [Uncaptioned image] (Guesmi et al., 2022) [Uncaptioned image] (Kouki et al., 2020) [Uncaptioned image] (Chatti et al., 2022) [Uncaptioned image], (Guesmi et al., 2022) [Uncaptioned image] (Chatti et al., 2022) [Uncaptioned image], (Guesmi et al., 2022) [Uncaptioned image] (Chatti et al., 2022) [Uncaptioned image], (Guesmi et al., 2022) [Uncaptioned image] (Chatti et al., 2022) [Uncaptioned image], (Guesmi et al., 2022) [Uncaptioned image]
Visualization Literacy (Guesmi et al., 2022) [Uncaptioned image] (Guesmi et al., 2022) [Uncaptioned image] (Hernandez-Bocanegra and Ziegler, 2020) [Uncaptioned image],(Kouki et al., 2020) [Uncaptioned image] (Guesmi et al., 2022) [Uncaptioned image] (Guesmi et al., 2022) [Uncaptioned image] (Guesmi et al., 2022) [Uncaptioned image], (Hernandez-Bocanegra and Ziegler, 2021) [Uncaptioned image] 18
$\times$ (Chatti et al., 2022) [Uncaptioned image] (Chatti et al., 2022) [Uncaptioned image] (Kouki et al., 2020) [Uncaptioned image] (Chatti et al., 2022) [Uncaptioned image], (Guesmi et al., 2022) [Uncaptioned image] (Chatti et al., 2022) [Uncaptioned image] (Chatti et al., 2022) [Uncaptioned image] (Chatti et al., 2022) [Uncaptioned image], (Guesmi et al., 2022) [Uncaptioned image]
Total (number of results) 7 4 21 9 13 12 10
$\times$ 8 8 18 18 13 12 14
Along with the explanation type and generation method, we investigated the results presented in Table 4 in the context of the application domain in which the explanations were evaluated (Table 5) and the type of recommender system that was explained (Table 6). We identified a total of nine different application domains that were evaluated, with point-of-interest recommendations being the most common (44 findings), followed by document (39 findings), and social recommendations (27 findings). We notice the dominance of some domains in certain user characteristics. The decision-making type and social awareness, for example, were almost exclusively evaluated in the point-of-interest domain and trust propensity and the education level of participants in the document and education domain, respectively.
Table 6. Number of results and publications in which the effect of user characteristics on explanation effects was analyzed and the type of recommender system as context variable. Rows with ✓indicate that an impact of the user characteristic was found; rows with $\times$ indicate that no impact was found. The last column sums the number of findings for each user characteristic. The last row sums up the findings where an effect was found or not found. The recommender type is shown after the domain and indicated by cbcb for content-based, cfcf for collaborative filtering, and hyhy for hybrid systems. Articles where information on any of the categories was not provided or the recommendations were simulated are marked with -.
User
Characteristic
Effect
Found
Effectiveness Efficiency
Perceived Explanation
Quality
Persuasiveness Transparency Trust Usability/UX Total
Demographic Age (Wilkinson et al., 2021)$cb$ (Nelekar et al., 2021)- 2
$\times$
Location (Berkovsky et al., 2018)- (Berkovsky et al., 2018)- 6
$\times$
Gender (Kleinerman et al., 2018)$cb$ (Ma et al., 2021)$cf$ (Ma et al., 2021)$cf$ (Ma et al., 2021)$cf$, (Tran et al., 2020)- 9
$\times$ (Wilkinson et al., 2021)$cb$, (Coba et al., 2019)- (Nelekar et al., 2021)-
Education (Zheng and Toribio, 2021)$hy$ (Ma et al., 2021)$cf$ 6
$\times$ (Ma et al., 2021)$cf$, (Zheng and Toribio, 2021)$hy$ (Ma et al., 2021)$cf$
Personality Agreeableness (Najafian et al., 2021a)- 4
$\times$ (Najafian et al., 2021b)- (Abdulrahman et al., 2022)$cb$
Conscientious- ness (Kouki et al., 2020)$hy$ (Nelekar et al., 2021)-, (Abdulrahman et al., 2022)$cb$ 4
$\times$ (Abdulrahman et al., 2022)$cb$
Decision-Making Type (Hernandez-Bocanegra and Ziegler, 2021, 2020)$cb$ (Hernandez-Bocanegra and Ziegler, 2020)$cb$ (Hernandez-Bocanegra and Ziegler, 2021, 2020)$cb$, (Coba et al., 2019)- (Hernandez-Bocanegra and Ziegler, 2021)$cb$ (Hernandez-Bocanegra and Ziegler, 2021, 2020)$cb$ (Hernandez-Bocanegra and Ziegler, 2021)$cb$ 17
$\times$ (Coba et al., 2019)- (Hernandez-Bocanegra and Ziegler, 2020)$cb$
Extraversion (Nelekar et al., 2021)- 4
$\times$ (Najafian et al., 2021b)- (Abdulrahman et al., 2022)$cb$
Need for Cognition (Martijn et al., 2022)$hy$ (Gajos and Mamykina, 2022)- 17
$\times$ (Chatti et al., 2022; Guesmi et al., 2022)$cb$ (Chatti et al., 2022; Guesmi et al., 2022)$cb$ (Liao and Sundar, 2021)cbcfcbcf, (Millecamp et al., 2020)- (Chatti et al., 2022; Guesmi et al., 2022)$cb$ (Gajos and Mamykina, 2022)-, (Chatti et al., 2022; Guesmi et al., 2022)$cb$ (Chatti et al., 2022; Guesmi et al., 2022)$cb$ (Chatti et al., 2022; Guesmi et al., 2022)$cb$
Neuroticism (Nelekar et al., 2021)-,(Kouki et al., 2020)$hy$ 7
$\times$ (Najafian et al., 2021a, b)- (Abdulrahman et al., 2022)$cb$
Openness (Kouki et al., 2020)$hy$ 3
$\times$ (Abdulrahman et al., 2022)$cb$
Personal Innovativeness (Chatti et al., 2022)$cb$ (Chatti et al., 2022)$cb$ (Guesmi et al., 2022)$cb$ (Chatti et al., 2022)$cb$ 13
$\times$ (Guesmi et al., 2022)$cb$ (Guesmi et al., 2022)$cb$ (Chatti et al., 2022; Guesmi et al., 2022)$cb$ (Chatti et al., 2022)$cb$ (Chatti et al., 2022; Guesmi et al., 2022)$cb$ (Chatti et al., 2022; Guesmi et al., 2022)$cb$
Trust Propensity (Nelekar et al., 2021)- (Guesmi et al., 2022)$cb$ 8
$\times$ (Guesmi et al., 2022)$cb$ (Guesmi et al., 2022)$cb$ (Guesmi et al., 2022)$cb$ (Guesmi et al., 2022)$cb$ (Guesmi et al., 2022)$cb$ (Guesmi et al., 2022)$cb$
Rationality 3
$\times$ (Hernandez-Bocanegra et al., 2020)- (Naveed et al., 2018)-
Social Awareness (Hernandez-Bocanegra and Ziegler, 2021, 2020)$cb$ (Hernandez-Bocanegra and Ziegler, 2021, 2020)$cb$ (Hernandez-Bocanegra and Ziegler, 2021)$cb$ (Starke et al., 2021b)-, (Hernandez-Bocanegra and Ziegler, 2021, 2020)$cb$ (Hernandez-Bocanegra et al., 2020)-, (Hernandez-Bocanegra and Ziegler, 2021, 2020)$cb$ 15
$\times$ (Hernandez-Bocanegra and Ziegler, 2020)$cb$ (Hernandez-Bocanegra et al., 2020)-
Trust in Technology (Loepp et al., 2019)$cf$ 1
$\times$
Valence (Kim et al., 2020)$cb$ 1
$\times$
Experience Domain Experience (Loepp et al., 2019)$cf$ 11
$\times$ (Guesmi et al., 2022)$cb$ (Guesmi et al., 2022)$cb$ (Kouki et al., 2020)$hy$ (Guesmi et al., 2022)$cb$, (Kouki et al., 2020)$hy$ (Guesmi et al., 2022)$cb$ (Guesmi et al., 2022)$cb$ (Guesmi et al., 2022)$cb$
Technical Expertise (Jacobs et al., 2021)- (Chatti et al., 2022)$cb$ (Kouki et al., 2020)$hy$ 16
$\times$ (Chatti et al., 2022; Guesmi et al., 2022)$cb$ (Guesmi et al., 2022)$cb$ (Kouki et al., 2020)$hy$ (Chatti et al., 2022; Guesmi et al., 2022)$cb$ (Chatti et al., 2022; Guesmi et al., 2022)$cb$ (Chatti et al., 2022; Guesmi et al., 2022)$cb$ (Chatti et al., 2022; Guesmi et al., 2022)$cb$
Visualization Literacy (Guesmi et al., 2022)$cb$ (Guesmi et al., 2022)$cb$ (Hernandez-Bocanegra and Ziegler, 2020)$cb$ ,(Kouki et al., 2020)$hy$ (Guesmi et al., 2022)$cb$ (Guesmi et al., 2022)$cb$ (Guesmi et al., 2022; Hernandez-Bocanegra and Ziegler, 2021)$cb$ 18
$\times$ (Chatti et al., 2022)$cb$ (Chatti et al., 2022)$cb$ (Kouki et al., 2020)$hy$ (Chatti et al., 2022; Guesmi et al., 2022)$cb$ (Chatti et al., 2022)$cb$ (Chatti et al., 2022)$cb$ (Chatti et al., 2022; Guesmi et al., 2022)$cb$
Total (number of results) 7 4 21 9 13 12 10
$\times$ 8 8 18 18 13 12 14
The dominant type of recommender system is content-based, with 107 out of 165 findings. Pure collaborative filtering systems are only used to generate the recommendations for 10 findings.